An H2SO4 solution of 13% by mass has density 1.09g mole inverse.its molarity and normality are

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
. Thickness of H2SO4 is 1.09 g/ml. 
wt of 1000ml=1lt=1.09*1000=1090 gms 
wt of H2SO4=1090*(13/100)=141.7 gms=141.7/98 moles=1.445moles 
molarity of H2SO4 arrangement is 1.445 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)